# Determinantal random subgraphs

### v2

@@ 15.4. Scaling limits
There is a scaling limit for the uniform spanning tree in fine mesh approximations of simply connected domains of the plane, given by the Schramm–Loewner evolution [LSW04].
In higher genus, does the random $g$-quasi-tree defined in Section 9.2 have a scaling limit, as suggested by the simulations in Figure 4 and 5? This question already arose from the work of [Lyo09].
The first case to consider would be the case of genus $g=1$. In that case, simple closed curves correspond to primitive homology classes $(p,q)\in H_{1}(\mathsf{T}^{2},\mathbb{Z})\simeq\mathbb{Z}^{2}$ with $\mathrm{gcd}(p,q)=1$. Can we describe the distribution on primitive homology classes induced by the simple curve in the case $\mathcal{C}_{1}$? Does it have a universal limit in the scaling limit (depending only on the torus’ aspect ratio)?
We could also consider the case of a discretization of a curved disk by a fine graph: for example, take a portion of the round sphere, discretize the $1$-form $dF$, where $F$ is the curvature $2$-form and consider the associated element of $\mathcal{C}_{1}(\mathsf{G})$. Does this random process have a scaling limit?
Before looking at two-dimensional scaling limits, one could also consider the case of a metric graph. For instance, we could also consider the Laplacian spanning forest on a cycle graph with $n$ edges. Does the process, or the process associated with the kernel $\mathsf{k}^{a}$ for $10<a<$ 1 (chosen appropriately with $n$), have a scaling limit on the unit circle when $n$ tends to infinity?

### v3

@@ 15.4. Scaling limits
The uniform spanning tree (UST) on fine mesh approximations of simply connected planar domains has a well-understood scaling limit described by the Schramm–Loewner evolution (SLE) [LSW04]. Do higher-genus analogues, such as $k$-quasitrees (for $k\leqslant 2g$), admit scaling limits as well? Simulations such as those in Figures 4 and 5 (which show uniform $2g$-quasitrees, which are Poincaré duals to uniform spanning trees) suggest this possibility, and this natural question already arises from [Lyo09].
Motivated by the construction in Section 12.1.1, one may ask whether random $g$-quasitrees associated with a distinguished Lagrangian subspace – chosen to ensure self-duality under Poincaré duality and conjectured criticality – also admit a scaling limit. Since the corresponding partition function can be expressed as a deformation of the spanning tree partition function by the determinant of the discrete normalized period matrix, and that the latter converges to its continuous counterpart under circle pattern approximations [Mer07, BS16], this suggests that proving scaling limit might be accessible when the Laplacian determinant is known to converge upon rescaling.
In genus $g=1$, simple closed curves correspond to primitive homology classes $(p,q)\in H_{1}(\mathsf{T}^{2},\mathbb{Z})\simeq\mathbb{Z}^{2}$ with $\gcd(p,q)=1$. Can we characterize the distribution on primitive classes induced by the random simple loop obtained from the determinantal $g$-quasitree? Does this distribution admit a universal scaling limit, depending only on the modulus of the torus?
We could also consider the case of a discretization of a curved disk by a fine graph: for example, take a portion of the round sphere, discretize the $1$-form $dF$, where $F$ is the curvature $2$-form and consider the determinantal random subgraph in $\mathcal{C}_{1}(\mathsf{G})$ defined by $\mathop{\mathrm{im}}d\oplus\mathbb{K}F$. Does this random spanning subgraph, or the unique random simple loop it defines, have a scaling limit?
Before tackling two-dimensional scaling limits, one could also consider simpler one-dimensional limits. For instance, we could consider the Laplacian spanning forest on a cycle graph with $n$ edges. Does the process associated with the kernel $\mathsf{k}$, or more likely, the process associated with the kernel $\mathsf{k}^{a}$ for $10<a<$ 1 (where $a$ is chosen appropriately with $n$), have a scaling limit on the unit circle when $n$ tends to infinity?